OpenStudy (anonymous):

What is the slope of the line which passes through (2, 5) and (3, 6)?

hero (hero):

1. Write down the slope-intercept formula: y = mx + b 2. Insert each point into the formula one at a time, creating two equations: 5 = 2m + b 6 = 3m + b 3. Isolate b in both equations: b = 5 - 2m b = 6 - 3m 4. Set b = b, then solve for m: 5 - 2m = 6 - 3m 3m - 2m = 6 - 5 (3 - 2)m = 1 m = 1 5. Thus the slope, m, between two points (2,5) and (3,6) is one.